OpenStudy (narissa):

Troy ate 3 more slices than Lisa together they ate 7 Lisa ate? domain of solution (1,2,3,4) a)define variables b)write algebraic expression each person ate

OpenStudy (photon336):

Let's say lisa ate n slices, well troy ate 3 more slices so it would be 3+n. Troy 3+n Lisa n combine 3+n+n = 3+2n =7 3+2n = 7 -3 -3 2n = 4 n = 2 so lisa ate 2 slices and troy ate 2+3 = 5 slices.

OpenStudy (photon336):

@narissa

OpenStudy (narissa):

c)write an equation to help solve d) solve the equation and write answer in complete sentence e) check answer

OpenStudy (narissa):

thank u u r smart and helpful

OpenStudy (photon336):

absolutely anytime @narissa do you get what I did though?

OpenStudy (photon336):

I'll give you another problem

OpenStudy (narissa):

yes can u help me with more stuff trying to get caught up

OpenStudy (photon336):

yeah no problem let's say jeff has two more CDs than joyce. they have a total of 8 cDs how many do each have?

OpenStudy (narissa):

````````````````````````thinking hold on

OpenStudy (photon336):

show me each step

OpenStudy (narissa):

2+n+n

OpenStudy (narissa):

k its hard

OpenStudy (narissa):

2+ or something

OpenStudy (narissa):

sorry im in a special class for math its hard for me

OpenStudy (photon336):

nice work @narissa

OpenStudy (photon336):

2+2n = 8

OpenStudy (photon336):

Now @narissa solve for n

OpenStudy (narissa):

4

OpenStudy (narissa):

oops 3

OpenStudy (photon336):

=)

OpenStudy (narissa):

2*3=6+2=8

OpenStudy (photon336):

so now we've found that 3 is the answer so how many do each have?

OpenStudy (narissa):

idk im confused sorry

OpenStudy (photon336):

dont worry so you've found that the answer is 3 here's the original question let's say jeff has two more CDs than joyce. they have a total of 8 cDs how many do each have?

OpenStudy (narissa):

jose has 6

OpenStudy (photon336):

Remember this we got n = 3 right? n+2 n 3+2 = 5 3 3+5 = 8 take a look at this what do you see>

OpenStudy (narissa):

oh jose has 5 and the other has 3

OpenStudy (photon336):

=D

OpenStudy (photon336):

that's basically the purpose behind what we've done

OpenStudy (narissa):

thanks for ur time

OpenStudy (narissa):

thank u

OpenStudy (narissa):

do u know what -4+5 is

OpenStudy (narissa):

is it 9

OpenStudy (photon336):

-4+5 = 1 why? is this the answer

OpenStudy (narissa):

no im going to post a new question k

OpenStudy (photon336):

ok